if the perimeter is 90 and the length is 7 inches shorter than triple the width, what are the demensions?

Answer by jorel1380(3719)

If you're talking about a rectangle, then 2L+2W = P. Since L = 3W-7, then
2(3W-7)+2W=90
6W-14+2W=90
8W=104
W=13
L=32.